OpenStudy (ksaimouli):

f(x)=under the root x+3 what is the range and domain?

OpenStudy (anonymous):

do you mean f(x) = sqrt(x+3) ?

OpenStudy (ksaimouli):

determine the domaine and rage

OpenStudy (ksaimouli):

\[\sqrt{?}\]+3

OpenStudy (ksaimouli):

under the root of x+3

OpenStudy (ksaimouli):

the root will be for both

OpenStudy (anonymous):

domain is (-3,INF)

OpenStudy (anonymous):

range is [0,INF) sorry domain is [-3,INF) '[' means including the value

OpenStudy (ksaimouli):

how

OpenStudy (ksaimouli):

u got the ans

OpenStudy (ksaimouli):

plz explain in detaile

OpenStudy (anonymous):

well - there is no real number which is the square root of a negative - therefore the lowest possible value of x+ 3 is zero. x+ 3 is zero when x = -3

OpenStudy (ksaimouli):

then how u got -3,inf

OpenStudy (anonymous):

well the domain is all the possible values of x - -3 is lowest value but values above it go up to infinity

OpenStudy (anonymous):

the range is all the possible values of f(x) lowest possible is sqrt(-3+3) = 0 and highest is infinity

OpenStudy (anonymous):

[-3,INF) means all values of x from -3 to infinity

OpenStudy (ksaimouli):

ok thanks